t points A 2.80 kg steel gas can holds 15.0 L of gasoline when Full. What is the average density (in kg/m2) of the full gas can, taking into account the volume occupied by steel as well as by gasoline? kg/m Addtional Materials Readng
Answer #1

Give Data

mass of steel can = mc = 2.80 kg

Volume of gasoline = Vg = 15 L = 0.015 m3

density of gasoline = \rhog = 680 kg/m3

Total mass of gasoline and can = m = mc + \rhog Vg = 2.80 + 680 x 0.015 =13 kg

density of steel , \rhos = 7600 kg/m3

Volume of steel can = Vs = mass of can / \rhos = 2.8/ 7600 = 0.000368 m3

Total Volume = Vs + Volume of can = 0.000368 + 0.015 = 0.0154

total density = total mass / total volume = 13/ 0.0154 = 844.15 kg/m3